(A)What would be the calculated magnetic moment of [Ni(H2O)6]2+ ?
(b) What kind of metals are extracted by electrometallurgy? Give an example.
(c) In the later actinides, why does the +3 state become predominent?
(d) What is YAG and what are its uses?
1
Expert's answer
2017-10-23T05:25:23-0400
a) 2.8 MB
b) K , Na , Ca , Al , Ti
Al2O3 + Na2AlF6 ( 960–970°С)
Al2O3 = Al(3+) + AlO3(3-) - melt
Al(3+) +3e(-) = Al(0) - Kat
4AlO3(3-) -12e = 2Al2O3 + 3O2 - An
2Al2O3 = 4Al + 3O2
c)They lose 7s2 electrons and one by one or with 6d or with 5f shells
d) Yttrium Aluminium Garnet , used in Nd:YAG laser